ONE EIGHTH OF THE SQUARE ROOT OF 7 LESS THAN A NUMBER IS 2 WHAT IS THE NUMBER?
I NEED MY DIPLOMA BUT IM STUCK!!!
i need all the steps so i can do the rest of the questions that are similar to this one
thanks

Answer by Photonjohn(42)

(1/8)(x-7)^1/2 = 2
(x-7)^1/2 = 16
square both sides of the equation
x-7 = 256
x = 256+7
x = 263